# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v4

@@ 2.2 Incidental Parameter Problem
as $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$. While $\widehat{\mu}$ is not affected by the IPP, the variance estimator $\widehat{\sigma}$ contains two non-negligible bias terms, $-\sigma_{0}/T$ and $-\sigma_{0}/N$. These bias terms cause $\sqrt{NT}(\widehat{\sigma}-\sigma_{0})$ to be centered at $-\kappa\sigma_{0}-\sigma_{0}/\kappa\neq 0$.
3 Bias Correction and Asymptotic Theory
The bias expansion of $l(\theta,\widehat{\psi}(\theta))$ in Equation (2.14) and Example 3 above motivate our approach of correcting the likelihood $l(\theta,\widehat{\psi}(\theta))$ by adding $B(\theta)$. The corrected likelihood can be viewed as an approximation to the IPP-free infeasible likelihood $l(\theta,\psi(\theta))$, with approximation error $o(1/\sqrt{NT})$. In Theorem 1 below we formalize the expansion in Equation (2.14). Specifically, we show that $B(\theta)$ can be further decomposed into $B(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o(1/\sqrt{NT})$, in which $B_{\alpha}(\theta)$ appears because of the estimation errors of the individual-specific heterogeneities $\alpha_{i}^{0}$, while $B_{\gamma}(\theta)$ is present due to the estimation errors of the time-specific heterogeneities $\gamma_{t}^{0}$. The exact formulas for both terms are given below. Theorem 2 follows to establish the corrected likelihood as
$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta)$ (3.1)
where we denote $\widehat{l}(\theta):=l(\theta,\widehat{\psi}(\theta))$ for simplicity from now on. $\widehat{B}_{\alpha}(\theta)$ and $\widehat{B}_{\gamma}(\theta)$ are estimates of $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$, respectively. The analytical formulas for $\widehat{B}_{\alpha}(\theta)$ and $\widehat{B}_{\gamma}(\theta)$ are provided below. We then show that
$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),$
where we use $l(\theta):=l(\theta,\psi(\theta))$ for simplicity below. This implies that $L(\theta)$ does not suffer from the IPP. The corrected estimator of $\theta_{0}$, and the LR, LM, and Wald test statistics, may subsequently be derived from $L\left(\theta\right)$, instead of $\widehat{l}(\theta)$. Specifically, the corrected estimator of $\theta_{0}$ is
$\widehat{\theta}_{L}:=\arg\max_{\theta}L(\theta).$
Under mild regularity conditions, $\widehat{\theta}_{L}$ is asymptotically unbiased, and its asymptotic properties align with the standard results of the classical MLE without the nuisance parameters. Moreover, the test statistics constructed from the corrected likelihood such as the Wald, LR, and LM also follow the limiting $\chi^{2}$ distribution under the null hypothesis. These results are formally stated in Corollaries 1 and 2, respectively.
Remark 2 (computation)
The computation of $\widehat{\theta}_{L}$ involves a two-layer maximization but can be done efficiently. In the inner layer, $\widehat{\psi}(\theta)$ is obtained by maximizing $l(\theta,\psi)$, keeping $\theta$ fixed. This is large-scale maximization problem and a few strategies can be exploited to expedite the procedure:
  1. 1.
    For many commonly used models, the first and second derivatives of $l(\theta,\psi)$ have clear analytical expressions, so that the standard Newton-Raphson algorithm can be employed to obtain $\widehat{\psi}(\theta)$.
  2. 2.
    Many models (specifically, generalized linear models) are equipped with linear indices, as in Example 1. In that case, the chain-rule-based technique in Sun and Dhaene (2019) can be used to achieve a fast calculation of the derivatives. The computational complexity of such a technique only grows with the number of linear indices, not the dimension of $\psi$, even without analytical expressions for the derivatives of $l(\theta,\psi)$.
  3. 3.
    The derivatives of the reparameterized $l(\theta,\psi)$ can be identically constructed from the derivatives of the original likelihood $\ell(\theta,\phi)$. Specifically, the reparameterization implies
    holding true for every $\phi=D^{\prime}\psi$ and every $\theta$.
In the outer layer where $L(\theta)$ is maximized w.r.t. $\theta$, it can be tedious to obtain analytical formulas for the derivatives of the corrected likelihood $L(\theta)$. However, the dimension of $\theta$ is usually small, so that invoking the Newton-Raphson algorithm with numerical derivatives generally does not suffer from too much computational pressure, provided that the inner-layer maximization is done efficiently.
@@ 3 Bias Correction and Asymptotic Theory
Now, we proceed to our bias correction procedure on the likelihood and present the related asymptotic theory. Denote $\widehat{\phi}(\theta):=D^{\prime}\widehat{\psi}(\theta)$ and $\phi(\theta):=D^{\prime}\psi(\theta)$, which are essentially the “unparameterized” counterparts of $\widehat{\psi}(\theta)$ and $\psi(\theta)$, respectively. We impose the following assumption.
@@
    $\sup_{N,T}\max_{1\leq i\leq N,1\leq t\leq T}\mathbb{E}\{[g(w_{it})]^{\eta}\}<\infty,$
    for some $\eta>2$.
@@
  • (vi) The matrix has $\sqrt{NT}\mathbb{E}H(\theta)$ has eigenvalues $\lambda_{p}(\theta)$ for $p=1,\ldots,(N+T-2)K$ which satisfy
@@
Here Assumption 1 is the same to Fernández-Val and Weidner (2016). It is used to ensure that the bias $\sqrt{NT}B(\theta)$ does not explode as $N,T\rightarrow\infty$, as discussed above. Assumption 1 is standard in the literature. See, e.g., Hahn and Newey (2004), Hahn and Kuersteiner (2011), and Arellano and Hahn (2016). It simply requires that the likelihood function $l_{i,t}(\theta,\phi)$ is differentiable in $\phi$ at least three times. The term $B(\theta)$ in (2.14) depends on the first and second derivatives, and the remainder $r(\theta)$ depends on the third. We require $l_{i,t}(\theta,\phi)$ and its derivatives to be uniformly bounded by a function $g(w_{it})$ whose moment exists to an order greater than the second. This is used to ensure, amongst other things, i) the interchangeability of $\mathbb{E}$ and $\partial$; and ii) the existence of moments, such as $\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}$ in (2.14), for every $\theta\in\Theta$. It is also used to establish the uniform rate of convergence, over $\theta\in\Theta$, of the corrected likelihood $L(\theta)$. The uniform convergence of $L(\theta)$ guarantees that the bias correction effect carries over to both the estimators and likelihood-based test statistics. Assumption 1 states that the true effects $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ sums up to 00 over $i=1,\ldots,N$ and $t=1,\ldots,T$. Given the additive structure of $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ in (2.1), this condition is without loss of generality because we merely let $\theta_{0}$ capture the average effects of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$. Assumption 1 is an identification condition for $\psi(\theta)$. It is standard for maximum likelihood models. See, e.g., Newey and McFadden (1994). Assumption 1 is common in the field of dynamic panels. See, for instance, Hahn and Kuersteiner (2007), Fernández-Val and Weidner (2016), and Dhaene and Jochmans (2015). Notice that, the argument is conditional on the true heterogeneous effects $\phi^{0}$. Without conditioning, $w_{it}$ may not be, for instance, independent over $i$, because $w_{it}$ and $w_{jt}$ may depend on the same $\gamma_{t}^{0}$. In addition, we only require the mixing coefficient $a_{i}(m)$ to decay at a polynomial rate such that $[a_{i}(m)]^{1-2/\eta}$ is summable as $m\rightarrow\infty$. The mixing condition is used to guarantee that certain autocovariances which appear in our proof are summable by Davydov’s inequality. Notice that we are not requiring $w_{it}$ to be identically distributed over $i$ or stationary over $t$. These are generally not possible because $w_{it}$ depends on $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$. Finally, Assumption 1 simply requires that the normalized expected Hessian $\sqrt{NT}\mathbb{E}H(\theta)$ has eigenvalues $\lambda_{p}(\theta)$ that are strictly less than 00 for all $N>N_{0}$ and $T>T_{0}$. This ensures that $\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}$ is well defined for large $N$ and $T$. This is a mild condition in that we only impose it on the expected Hessian $\mathbb{E}H(\theta)$ of the reparameterized likelihood. We state the condition with “less than 00” because the Hessian in a maximum likelihood model is usually negative definite.
Next, partition [^5] Note that $s(\theta)=s(\theta)-\mathbb{E}s\left(\theta\right)$ because $\mathbb{E}s\left(\theta\right)=0$ by definition. We keep $\mathbb{E}s\left(\theta\right)$ here for the convenience of constructing the corrected likelihood. $s(\theta)-\mathbb{E}s\left(\theta\right)$ and $[\mathbb{E}H(\theta)]^{-1}$ as
@@
$s(\theta)-\mathbb{E}s\left(\theta\right)=:\left(\begin{array}[c]{c}\widetilde{s}_{\alpha}(\theta)\\ \widetilde{s}_{\gamma}(\theta)\end{array}\right),\qquad[\mathbb{E}H(\theta)]^{-1}=:\left(\begin{array}[c]{cc}H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)&H_{\alpha\gamma}^{\mathcal{\ast}}(\theta)\\ H_{\gamma\alpha}^{\mathcal{\ast}}(\theta)&H_{\gamma\gamma}^{\mathcal{\ast}}(\theta)\end{array}\right),$ (3.2)
where $\widetilde{s}_{\alpha}(\theta)$ is $(N-1)K\times 1$ and $\widetilde{s}_{\gamma}(\theta)$ is $(T-1)K\times 1$ while $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ is $(N-1)K\times(N-1)K$ and $H_{\gamma\gamma}^{\ast}(\theta)$ is $(T-1)K\times(T-1)K$. The off-diagonal blocks $H_{\alpha\gamma}^{\ast}(\theta)$ and $H_{\gamma\alpha}^{\mathcal{\ast}}(\theta)$ are defined accordingly. Further denote
$S_{\alpha\alpha}(\theta):=\mathbb{E\{}\widetilde{s}_{\alpha}(\theta)\mathbb{[}\widetilde{s}_{\alpha}(\theta)]^{\prime}\},\qquad S_{\gamma\gamma}(\theta):=\mathbb{E\{}\widetilde{s}_{\gamma}(\theta)\mathbb{[}\widetilde{s}_{\gamma}(\theta)]^{\prime}\}.$
Theorem 1
Under Assumption 1 and uniformly over $\theta\in\Theta$,
as $N,T\rightarrow\infty$, where
@@
$B_{\alpha}(\theta):=\frac{1}{2}\operatorname*{trace}[S_{\alpha\alpha}(\theta)H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad B_{\gamma}(\theta):=\frac{1}{2}\operatorname*{trace}\mathbb{[}S_{\gamma\gamma}(\theta)H_{\gamma\gamma}^{\ast}(\theta)],$
and $B_{\alpha}(\theta)+B_{\gamma}(\theta)=O(1/\sqrt{NT})$.
[…]
The bias terms $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$ involves $\psi(\theta)$ and $\mathbb{E}$ which are infeasible to compute. However, $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$ can be estimated by plug-in estimates of $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$, $H_{\gamma\gamma}^{\ast}(\theta)$, $S_{\alpha\alpha}(\theta)$, and $S_{\gamma\gamma}(\theta)$. The two Hessian terms can be estimated in a straightforward way. Denote $\widehat{H}(\theta):=H(\theta,\widehat{\psi}(\theta))$ and partition $[\widehat{H}(\theta)]^{-1}$ into

### v5

@@ A Examples
as $N,T\rightarrow\infty$ with $N/T\rightarrow\kappa$. While $\widehat{\mu}$ is not affected by the IPP, the variance estimator $\widehat{\sigma}$ contains two non-negligible bias terms, $\sigma_{0}/T$ and $\sigma_{0}/N$. These bias terms cause $\sqrt{NT}(\widehat{\sigma}-\sigma_{0})$ to be centered at $\kappa\sigma_{0}-\sigma_{0}/\kappa\neq 0$.
B Remarks
Remark 4 (uniform inference)
The inference based on our corrected likelihood is uniformly valid regardless of whether the underlying data-generating process has a homogeneous or heterogeneous slope coefficient. This is similar to Lu and Su (2023), except that they consider a linear regression. For instance, consider the data-generating process
$Y_{it}=\mathds{1}\{\left(\beta_{0}+\alpha_{1,i}^{0}+\gamma_{1,t}^{0}\right)X_{it}+\alpha_{2,i}^{0}+\gamma_{2,t}^{0}+\varepsilon_{it}>0\},$
where the slope coefficient of $X_{it}$ can be homogeneous, if $\alpha_{1,i}^{0}=\gamma_{1,t}^{0}=0$ for all $i$ and $t$, or heterogeneous, if otherwise. In either case, inference based on the corrected likelihood under the model
$Y_{it}=\mathds{1}\{(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\},\ \alpha_{1,i}\neq 0\ \text{and }\gamma_{1,t}\neq 0,$
remains valid. This uniform validity shares a similar spirit as the uniform inference in Lu and Su (2023). In fact, the homogeneous data-generating process is just a special case of the heterogeneous one. By employing the bias-corrected likelihood for the heterogeneous model, inference on $\beta$ remains valid, albeit with a possible loss of efficiency when the true data-generating process is homogeneous.
In addition, we treat the heterogeneous variables $\alpha_{k,i}^{0}$ and $\gamma_{k,t}^{0}$ as parameters by conditioning on them, which is common for panel maximum likelihood methods, we do not further distinguish between locally heterogeneous and fully heterogeneous cases as in Lu and Su (2023).
Remark 5 (expectation)
Remark 6 (truncation)
In the estimator $\widehat{S}_{\alpha\alpha}(\theta)$, we use the truncation technique commonly found in e.g., Hahn and Kuersteiner (2011) (see also Hahn and Kuersteiner (2007)), Fernández-Val and Weidner (2016), and Galvao and Kato (2016). This is related to the literature on consistent estimations of autocovariances, $\mathbb{E\{}\widetilde{s}_{i,t}^{\alpha}(\theta)[\widetilde{s}_{i,s}^{\alpha}(\theta)]\}$ in our case. Detailed discussions in this regard can be found in Andrews (1991). Without truncation, we would have $\widehat{S}_{\alpha\alpha}(\theta)=\widehat{s}_{\alpha}(\theta)[\widehat{s}_{\alpha}(\theta)]^{\prime}=0$ for every $\theta$ because $\widehat{s}_{\alpha}(\theta)=0$ by definition, where $\widehat{s}_{\alpha}(\theta)$ is defined from $\widehat{s}(\theta):=s(\theta,\widehat{\psi}(\theta))$ under the same partition in (3.2). Here, the truncation parameter $\tau\rightarrow\infty$ but sufficiently slower relative to $T$ (see Theorem 2). Hahn and Kuersteiner (2011) suggests the default value of $\tau=1$ for small $T$. For intermediate $T$, we recommend following the advice from Fernández-Val and Weidner (2016) to conduct a sensitivity test for different values of $\tau$ and to avoid using very large $\tau$ values. In our simulation, we use $\tau=1$ and $2$ and find the difference relatively insignificant.
On the other hand, the term $\widehat{S}_{\gamma\gamma}(\theta)$ does not require any truncation, because we have already eliminated $\mathbb{E\{}\widetilde{s}_{i}^{\gamma}(\theta)[\widetilde{s}_{j}^{\gamma}(\theta)]^{\prime}\}$ for $i\neq j$ from Equation (C.1) using cross-sectional independence. One may still opt to truncate $\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$, such that all elements beyond the $\tau$-th sub- and super-diagonal are symmetrically set to 0. We have experimented with such truncations but found little impact on $\widehat{\theta}_{L}$ and test statistics. The comparison is, hence, not reported in the paper.
C Technical Details
C.1 Bias Terms
The reparameterization ensures the following identity:
where $\mathcal{S}_{\alpha\alpha}(\theta)$ is the outer product of the $NK\times 1$ vector of $\partial_{\alpha}\ell(\theta,\phi)-\mathbb{E}\partial_{\alpha}\ell(\theta,\phi)$ evaluated at $\phi=\phi(\theta)$, where $\partial_{\alpha}\ell(\theta,\phi)$ is the derivative of the original likelihood $\ell(\theta,\phi)$ w.r.t. $\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$. Specifically,
$\widetilde{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\alpha}(\theta,\phi(\theta)).$
Under cross-sectional independence, $\mathcal{S}_{\alpha\alpha}(\theta)$ simplifies to
Intuitively, $\widehat{S}_{\alpha\alpha}(\theta)$ in (3.6) is a natural plug-in estimator of $S_{\alpha\alpha}(\theta)$.
Similarly for $S_{\gamma\gamma}(\theta)$, the reparameterization produces the identity:
$\widetilde{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\gamma}(\theta,\phi(\theta)).$
By cross-sectional independence, $\mathcal{S}_{\gamma\gamma}(\theta)$ similarly reduces to
$\mathcal{S}_{\gamma\gamma}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\mathbb{E\{}s_{i}^{\gamma}(\theta)[s_{i}^{\gamma}(\theta)]^{\prime}\}.$ (C.1)
Naturally, $\widehat{S}_{\gamma\gamma}(\theta)$ in (3.7) is also an intuitive plug-in estimator of $S_{\gamma\gamma}(\theta)$.
C.2 Regularity Conditions
@@
Denote $\phi(\theta):=D^{\prime}\psi(\theta)$, which is the “unparameterized” counterpart of $\psi(\theta)$. We impose the following assumption to establish the asymptotic theory for the expansion in (3.1) and for the corrected likelihood $L(\theta)$.
@@
    $\sup_{N,T}\max_{1\leq i\leq N,1\leq t\leq T}\mathbb{E}_{\phi}\{[g(w_{it})]^{\eta}\}<\infty$
    for some $\eta>2,$ where $\mathbb{E}_{\phi}$ denotes the conditional expectation w.r.t. the joint distribution of $w_{it}$, given the heterogeneous effects $\phi^{0}$.
@@
  • (vi) The matrix $\sqrt{NT}\mathbb{E}[H(\theta)]$ has eigenvalues $\lambda_{p}(\theta)$ for $p=1,\ldots,(N+T-2)K$ which satisfy
@@
Assumption 1 is the same to Fernández-Val and Weidner (2016). It is used to ensure that the bias $\sqrt{NT}B(\theta)$ does not explode as $N,T\rightarrow\infty$, as discussed above. Assumption 1 is standard in the literature. See, e.g., Hahn and Newey (2004), Hahn and Kuersteiner (2011), and Arellano and Hahn (2016). It simply requires that the likelihood function $l_{i,t}(\theta,\phi)$ is differentiable in $\phi$ at least three times. The term $B(\theta)$ in (2.7) depends on the first and second derivatives, and the remainder $r(\theta)$ depends on the third. We require $l_{i,t}(\theta,\phi)$ and its derivatives to be uniformly bounded by a function $g(w_{it})$ whose moment exists to an order greater than the second. This is used to ensure, amongst other things, i) the interchangeability of $\mathbb{E}$ and $\partial$; and ii) the existence of moments, such as $\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}$ in (2.7), for every $\theta\in\Theta$. It is also used to establish the uniform rate of convergence, over $\theta\in\Theta$, of the corrected likelihood $L(\theta)$. The uniform convergence of $L(\theta)$ guarantees that the bias correction effect carries over to both the estimators and likelihood-based test statistics. Assumption 1 states that the true effects $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$ sums up to 0 over $i=1,\ldots,N$ and $t=1,\ldots,T$. Given the additive structure of $\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}$ in (2.1), this condition is without loss of generality because we merely let $\theta_{0}$ capture the average effects of $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$. Assumption 1 is an identification condition for $\psi(\theta)$. It is standard for maximum likelihood models. See, e.g., Newey and McFadden (1994). Assumption 1 is common in the field of dynamic panels. See, for instance, Hahn and Kuersteiner (2007), Fernández-Val and Weidner (2016), and Dhaene and Jochmans (2015). Notice that, the argument is conditional on the true heterogeneous effects $\phi^{0}$. Without conditioning, $w_{it}$ may not be, for instance, independent over $i$, because $w_{it}$ and $w_{jt}$ may depend on the same $\gamma_{t}^{0}$. In addition, we only require the mixing coefficient $a_{i}(m)$ to decay at a polynomial rate such that $a_{i}(m)]^{1-2/\eta}$ is summable as $m\rightarrow\infty$. The mixing condition is used to guarantee that certain autocovariances which appear in our proof are summable by Davydov’s inequality. Notice that we are not requiring $w_{it}$ to be identically distributed over $i$ or stationary over $t$. These are generally not possible because $w_{it}$ depends on $\alpha_{i}^{0}$ and $\gamma_{t}^{0}$. Finally, Assumption 1 simply requires that the normalized expected Hessian $\sqrt{NT}\mathbb{E}H(\theta)$ has eigenvalues $\lambda_{p}(\theta)$ that are strictly less than 0 for all $N>N_{0}$ and $T>T_{0}$. This ensures that $\mathbb{E\{}s^{\prime}(\theta)[H(\theta)]^{-1}s(\theta)\}$ is well defined for large $N$ and $T$. This is a mild condition in that we only impose it on the expected Hessian $\mathbb{E}H(\theta)$ of the reparameterized likelihood, which does not suffer from the indeterminacy in Equation (2.3). We state the condition with “less than 0” because the Hessian in a maximum likelihood model is usually negative definite.
The following Lemma provides low-level sufficient conditions for Assumptions 1 and 1 under a logit model
$Y_{it}=\mathds{1}\{X_{it}^{\prime}(\theta+\alpha_{i}+\gamma_{t})+\varepsilon_{it}>0\},$ (C.2)
where $\varepsilon_{it}$ follows a standard logistic distribution independent of $X_{it}$, $\alpha_{i}$, and $\gamma_{t}.$
Lemma C.1
Under model (C.2), a sufficient condition of Assumption 1 is that $\sup_{i,t}\mathbb{E}_{\phi}\|X_{it}\|_{\max}^{3\eta}$$\infty$ for some $\eta>2.$ Let $X_{i\cdot}:=\left(X_{i1},\ldots,X_{iT}\right)^{\prime}$ and $X_{\cdot t}:=\left(X_{1t},\ldots,X_{Nt}\right)^{\prime}$. Assumption 1 is satisfied if, for all $i$, the columns of $X_{i\cdot}$ are linearly independent and, for all $t$, the columns of $X_{\cdot t}$ are linearly independent.
Proof. Under the logit model (C.2) the log-likelihood for individual $i$ at time $t$ is
$l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)=\left(Y_{it}-1\right)X_{it}^{\prime}(\theta+\alpha_{i}+\gamma_{t})-\log\big(1+\exp\left(-X_{it}^{\prime}(\theta+\alpha_{i}+\gamma_{t})\right)\big).$
We assume that $X_{it}$ is scalar regressor for simplicity. The derivatives of $l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)$ w.r.t. the heterogeneous parameters are, without loss of generality,
$\partial_{\alpha_{i}}l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)=\left(Y_{it}-\frac{1}{1+\exp(-X_{it}(\theta+\alpha_{i}+\gamma_{t}))}\right)X_{it},$
$\partial_{\alpha_{i}\gamma_{t}}l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)=\frac{-\exp(-X_{it}(\theta+\alpha_{i}+\gamma_{t}))}{[1+\exp(-X_{it}(\theta+\alpha_{i}+\gamma_{t}))]^{2}}X_{it}^{2},$
and
$\partial_{\alpha_{i}\gamma_{t}\gamma_{t}}l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)=\frac{\exp(-X_{it}(\theta+\alpha_{i}+\gamma_{t}))\left[1-\exp(-X_{it}(\theta+\alpha_{i}+\gamma_{t}))\right]}{[1+\exp(-X_{it}(\theta+\alpha_{i}+\gamma_{t}))]^{3}}X_{it}^{3}.$
Note that the cross derivatives w.r.t. $\alpha_{i}$ and $\alpha_{j},$ or $\gamma_{t}$ and $\gamma_{s},$ are all 0 for $i\neq j$ or $t\neq s$.
Assumption 1. We have
$\sup_{\theta,\alpha_{i},\gamma_{t}}\left|l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)\right|$ $\leq\sup_{\theta,\alpha_{i},\gamma_{t}}\left|X_{it}(\theta+\alpha_{i}+\gamma_{t})\right|+\sup_{\theta,\alpha_{i},\gamma_{t}}\log\big(1+\exp\left(-X_{it}(\theta+\alpha_{i}+\gamma_{t})\right)\big)$
$\leq 2\sup_{\theta,\alpha_{i},\gamma_{t}}\left|X_{it}(\theta+\alpha_{i}+\gamma_{t})\right|+\log 2$
$\leq C|X_{it}|+\log 2,$ (C.3)
where $C<\infty$ is a positive constant depending on the compact parameter spaces $\Theta$ and $\Phi.$ It is easy to see that
$\sup_{\theta,\alpha_{i},\gamma_{t}}\left|\partial_{\alpha_{i}}l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)\right|\leq\left|X_{it}\right|.$ (C.4)
For the second derivative, observe that
@@
$\sup_{\theta,\alpha_{i},\gamma_{t}}\left|\partial_{\alpha_{i}\gamma_{t}}l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)\right|\leq\sup_{\pi\in\mathbb{R}}\frac{\exp(-\pi)}{(\exp(-\pi)+1)^{2}}X_{it}^{2}\leq\frac{1}{4}X_{it}^{2}.$
For the third derivative, we have
@@
$\sup_{\theta,\alpha_{i},\gamma_{t}}\left|\partial_{\alpha_{i}\gamma_{t}\gamma_{t}}l_{i,t}\left(\theta,\alpha_{i},\gamma_{t}\right)\right|$ $\leq\sup_{\pi\in\mathbb{R}}\left|\frac{\exp(-\pi)\left(1-\exp(-\pi)\right)}{(\exp(-\pi)+1)^{3}}\right|\left|X_{it}\right|^{3}$
$\leq\sup_{s\geq 0}\frac{s+s^{2}}{(1+s)^{3}}\left|X_{it}\right|^{3}$
$\leq\left(\frac{1}{5}+\frac{1}{3}\right)\left|X_{it}\right|^{3}.$ (C.5)
Combining (C.3)–(C.5), Assumption 1 holds with
$g\left(w_{it}\right)=C(\left\|X_{it}\right\|_{\max}+1)^{3}+\log 2,$
where $C>1.$ Therefore, the condition $\sup_{i,t}\mathbb{E}_{\phi}\left\|X_{it}\right\|_{\max}^{3\eta}<\infty$ is sufficient for Assumption 1.
Assumption 1. Under the logit model (C.2), the Hessian can be expressed as
$\sqrt{NT}H(\theta)=\frac{1}{\sqrt{NT}}DX^{\prime}H^{\pi}(\theta)XD^{\prime},$ (C.6)
where $D$ is the deterministic matrix defined in Section 2.1, and $X=(X^{\alpha},X^{\gamma})$ with